# Uncertainty quantification for data-driven weather models

### v1

@@ 1 Introduction
Modern weather forecasts are usually based on simulations from physics-based numerical weather prediction (NWP) models, which describe atmospheric processes via systems of partial differential equations. To quantify forecast uncertainty and provide probabilistic predictions, NWP models are typically run several times with varying initial conditions and perturbed model physics, resulting in an ensemble of predictions. Numerically solving the differential equations requires tremendous computational resources, limiting the spatial and temporal resolution, as well as the number of ensemble runs. The history of NWP since its inception around 70 years ago has been a success story, albeit a “quiet” one characterized by continued, small improvements through the steady accumulation of scientific knowledge and technological advances (Bauer et al., 2015).
@@
Currently, a major leap in the formerly quiet success story of NWP can be observed due to the unprecedented success and rapid advancement of purely data-driven machine learning (ML) models for weather prediction. Contrary to NWP, data-driven weather models do not include any physics-based equations and aim to predict the future weather state (typically iteratively in steps of hours to days) from the initial weather state only, using statistical relations learned from past data. Beyond improved forecasts, the major advantages of data-driven models are their substantially lower computational costs (and accompanied energy consumption) and the faster generation of forecasts, once a model has been trained. Over the past two years, fundamental advances have been achieved, with purely data-driven weather models now convincingly outperforming state-of-the-art NWP systems, as recently reviewed in Ben-Bouallègue et al. (2023a). The most notable contributions and models include Keisler (2022), FourCastNet (Pathak et al., 2022), Pangu-Weather (Bi et al., 2023), GraphCast (Lam et al., 2022), ClimaX (Nguyen et al., 2023a), FengWu (Chen et al., 2023a), FuXi (Chen et al., 2023c), SwinRDM (Chen et al., 2023b), AtmoRep (Lessig et al., 2023), NeuralGCM (Kochkov et al., 2023), Stormer (Nguyen et al., 2023b), and GenCast (Price et al., 2023). All models utilize the ERA5 global reanalysis dataset (Hersbach et al., 2020) for training and evaluation, and are run at grid spacings of up to ${}^{\circ}$.
@@
However, most of these efforts have been focused on deterministic forecasts only, making it impossible to quantify forecast uncertainties which is crucial for optimal decision making, and one of the reasons underlying a transdisciplinary transition towards probabilistic forecasts (Gneiting and Katzfuss, 2014). Therefore, the overarching aim of our work is to investigate approaches to generate probabilistic predictions from deterministic data-driven weather models. [^3] An ideal solution to this challenge might be inherently probabilistic data-driven approaches, for example generative ML methods. The AtmoRep (Lessig et al., 2023), NeuralGCM (Kochkov et al., 2023) and GenCast (Price et al., 2023) models represent first steps in this direction. In order to generate probabilistic forecasts, we consider two main approaches for uncertainty quantification (UQ) for data-driven weather models. A schematic overview is provided in Figure 1.
@@
Initial condition (IC)-based approaches generate an ensemble forecast by running a data-driven model multiple times based on a number of (slightly) different initial conditions. These initial condition ensembles can be generated in various ways, and we consider three variants: Adding random noise to the initial weather state obtained from the ERA5 dataset, as first proposed by Pathak et al. (2022); utilizing the perturbed initial conditions of a physics-based NWP ensemble model (Buizza et al., 2008); and generating conditions based on perturbations computed from randomly selected past data, as proposed by Magnusson et al. (2009). IC approaches generally require the capabilities to run the data-driven models for a set of input data (i.e., estimated models, code, data, and suitable hardware infrastructure), which has recently become possible since code and data have been made public for some of the models (most notably FourCastNet, Pangu-Weather, and GraphCast), but still poses technical challenges due to the substantial computing and disk space requirements.
Post-hoc (PH) UQ approaches, by contrast, utilize statistical or ML methods to supplement deterministic forecasts with uncertainty information and thus turn them into probabilistic forecasts. These methods only require a training dataset of deterministic forecasts and corresponding observations. A large variety of such approaches has been proposed, e.g., conformal prediction (Angelopoulos and Bates, 2021) or distributional regression (Gneiting and Katzfuss, 2014). Here, we consider two distributional regression approaches particularly relevant for atmospheric science applications, where such methods have been mostly used in the context of statistical post-processing to correct systematic errors of NWP forecasts (Vannitsem et al., 2021). The EasyUQ (Walz et al., 2024a) approach builds on the recent isotonic distributional regression technique (Henzi et al., 2021) and yields statistically optimal discrete predictive distributions by leveraging the pool-adjacent-violators algorithm for nonparametric isotonic regression. EasyUQ utilizes deterministic forecasts of the target variable as sole input, and has, e.g., recently been used in Walz et al. (2024b) to generate probabilistic forecasts of precipitation from deterministic inputs. Over the past years, modern ensemble post-processing methods based on neural networks have been proposed which enable the incorporation of additional input variables and the data-driven learning of complex relationships between the inputs and distribution forecasts. We will build on the parametric distributional regression network approach first proposed in Rasp and Lerch (2018), which has been successfully extended for many target variables (e.g., in Schulz and Lerch, 2022) and has been used to generate corrected probabilistic forecasts from deterministic inputs from an NWP model (Chapman et al., 2022; Gneiting et al., 2023).
@@ 2 Data and setup
Our study focuses on the Pangu-Weather model developed by Bi et al. (2023). Additional results for the FourCastNet model (Pathak et al., 2022) are available in the supplemental material. Pangu-Weather is a three-dimensional vision transformer architecture with specific adaptations and extensions to weather prediction, and was one of the first data-driven models to achieve improvements over physics-based NWP models. The Pangu-Weather model produces global forecasts of five atmospheric variables (Z, Q, T, U, V) on 13 pressure levels and four surface variables (MSL, U10M, V10M, T2M) at a grid spacing of ${}^{\circ}$. It is trained based on 39 years of ERA5 reanalysis data from 1979–2017. In Bi et al. (2023), data from the year 2018 was used as validation data, and data from 2019 serves as a test dataset. For details regarding the model architecture, training procedure, and forecast quality, we refer to Bi et al. (2023). To implement the UQ methods described below, we adapted Pangu-Weather code and data provided by the authors [^5] https://github.com/198808xc/Pangu-Weather for our purposes.
@@
As a reference forecast but also as initial condition perturbations (see Section 4, we retrieve operational ensemble forecasts of ECMWF’s ensemble prediction system, which is based on the ECMWF Integrated Forecasting System (IFS). The data are retrieved for the same training and evaluation periods on a regular latitude-longitude grid of 0.25$\times$0.25° covering the identical spatial domain and forecast lead times. It should be noted that the native spatial resolution of the ensemble prediction system is slightly higher than that of ERA5 which may cause differences in regions of high topography.
3 Notation
@@ 4.1 Initial condition ensemble approaches
The general idea behind all considered IC ensemble approaches is that based on (slightly) different initial conditions
@@
The IC approaches described below mainly differ in the way the IC ensemble $\bm{X}_{l,t,0}$ is generated. Specifically, we consider Gaussian noise perturbations, random field perturbations and IFS perturbed initial conditions, which are introduced in detail below. Exemplary perturbations for a common initialization date are visualized in Figure 2. A natural shortcoming of all IC approaches is that they are inherently limited to accounting for initial condition uncertainty only, and not for model uncertainty, which is, e.g., addressed in physics-based NWP models via stochastic parametrizations of subgrid processes (Palmer, 2019b).
@@ 4.1.1 Gaussian noise perturbations (GNP)
$X_{l,t,0}^{\text{GNP},m}=Y_{l,t}+\varepsilon_{l,t}\quad\text{for }m=1,...,M,$
where $\varepsilon_{l,t}\sim\mathcal{N}(0,\gamma\,\sigma)$, $\sigma$ denotes the mean standard deviation of the respective variable over all grid points, and $\gamma$ is a tuning parameter. Samples of the Gaussian noise process are thus generated independently over members $m$, locations $l$, variables and initialization times $t$. While Pathak et al. (2022) use $\gamma=0.3$ for the FourCastNet model, our initial experiments with the Pangu-Weather model suggested that $\gamma=0.001$ leads to better results.
@@
4.1.2 Random field perturbations (RFP)
An alternative, data-driven approach to generate IC ensembles, which we will refer to as random field perturbations, was proposed by Magnusson et al. (2009) in the context of physics-based NWP ensemble models. They argue that adding noise to the initial conditions ignores the underlying dynamics of the weather system. Instead, they suggest to use the scaled difference of two independent, randomly selected atmospheric states from the past as perturbation, which has the advantage of preserving linear balances in the system. The random field perturbations are calculated as
where $\bm{Y}_{\tau_{i}^{m}}=\{Y_{l,\tau_{i}^{m}},l\in\mathbb{L}_{G}\},i=1,2$ denotes the global observed ERA5 field of the variable of interest at date $\tau_{i}^{m}$, and $\|\cdot\|_{\mathrm{Etot}}$ denotes the total energy norm. We choose the dates $\tau_{1}^{m},\tau_{2}^{m}$ randomly from the same month as $t$ from the training dataset (2018–2021) to account for seasonal variability, but sample $\tau_{1}^{m}$ and $\tau_{2}^{m}$ from different years to ensure (approximate) independence. The constant $\alpha$ is a tuning parameter and controls the dispersion of the IC ensemble. Based on some preliminary tests, we chose $\alpha=5\cdot 10^{6}$ for our case study.
With these choices, global perturbations $\bm{\xi}_{t}^{m,\alpha},m=1,...,M$, are computed and added to the corresponding ERA5 initial conditions, i.e.,
$X_{l,t,0}^{\text{RFP},m}=Y_{l,t}+\xi^{m,\alpha}_{l,t}$
for $m=1,...,M$ and all $l\in\mathbb{L}_{G}$.
@@ 4.1.3 IFS initial conditions (IFSP)
Finally, we consider an IC approach more akin to the operational practice of running NWP ensemble models (Palmer, 2019a) by utilizing the (ensemble of) initial conditions of the ECMWF ensemble model to initialize the deterministic data-driven model. Specifically, we select the values at initialization time (i.e., the forecasts at step $h=0$) of the perturbed members $\bm{Z}_{t,0}^{m}=\{Z_{l,t,0}^{m},l\in\mathbb{L}_{G}\},m=1,...,M$, of the ECMWF ensemble, i.e.,
@@ 4.2 Post-hoc approaches
is available, where $t$ denotes an initialalization time in the training dataset (2018–2021).
Based on the training dataset, the PH methods yield forecast distributions $F_{l,t,h}$. Given a deterministic Pangu-Weather forecast $X^{\text{PW}}_{l,t^{\ast},h}$ in the test dataset (2022), a probabilistic forecast for the date $t^{\ast}$ and lead time $h$ can thus be obtained by using the corresponding deterministic forecast as input to the trained PH model. In the following, we consider two complementary PH methods based on statistical and ML approaches.
@@ 4.2.1 EasyUQ
EasyUQ, proposed by Walz et al. (2024a), aims at learning a predictive distribution from deterministic, single-valued model output. EasyUQ proceeds separately for every location $l\in\mathbb{L}_{E}$ and lead time $h$. To simplify notation, we will suppress the lead time index $h$ in the current subsection, and note that all forecasts and observations should be understood as those for the corresponding lead time only. Given corresponding data of the form $(X^{\text{PW}}_{l,t},Y_{l,t}),t=1,...,T$, and assuming that the predictive cumulative distribution functions (CDFs) $F_{x}(y)=\mathbb{P}(Y_{l,t}\leq y|X_{l,t}=x)$ are increasing in stochastic order in $x$, i.e., $F_{x}(y)\geq F_{x^{\prime}}(y)$ for all $y\in\mathbb{R}$ if $x\geq x^{\prime}$, the IDR-estimated predictive CDF is then given by
$\hat{F}^{\text{IDR}}_{l,t}(y)=\hat{F}^{\text{IDR}}_{X^{\text{PW}}_{l,t}}(y)=\min_{k=1,\ldots,t}\max_{\ell=t,\ldots,T}\frac{1}{\ell-k+1}\sum_{t^{\prime}=k}^{\ell}\mathbbm{1}\{Y_{l,t^{\prime}}\leq y\},\quad t=1,\ldots,T.$
Building on the recent isotonic distributional regression (IDR) technique (Henzi et al., 2021), $\hat{F}^{\text{IDR}}_{l,t}$ is a statistically optimal discrete predictive distribution in that it minimizes the continuous ranked probability score (CRPS, see Section 4.3) over all conditional distributions satisfying the assumption of stochastic ordering. For theoretical results and more details on EasyUQ and IDR, we refer to Walz et al. (2024a) and Henzi et al. (2021). EasyUQ does not require any choices of tuning parameters and thus constitutes an attractive benchmark method that can be applied in a fully automated manner. Walz et al. (2024a) note that EasyUQ yields similar forecast performance as conformal prediction, and in case studies on post-processing, EasyUQ showed predictive performance comparable to other statistical models (e.g., Schulz and Lerch, 2022).

### v2

@@ 1 Introduction
Modern weather forecasts are usually based on simulations from physics-based numerical weather prediction (NWP) models, which describe atmospheric processes via systems of partial differential equations. To quantify forecast uncertainty and provide probabilistic predictions, NWP models are typically run several times with varying initial conditions and perturbed model physics, resulting in an ensemble of predictions. Numerically solving the differential equations requires tremendous computational resources, limiting the spatial resolution, as well as the number of ensemble runs. The history of NWP since its inception around 70 years ago has been a success story, albeit a “quiet” one characterized by continued, small improvements through the steady accumulation of scientific knowledge and technological advances (Bauer et al., 2015).
@@
Currently, a major leap in the formerly quiet success story of NWP can be observed due to the unprecedented success and rapid advancement of purely data-driven machine learning (ML) models for weather prediction. Contrary to NWP, data-driven weather models do not include any physics-based equations and aim to predict the future weather state (typically iteratively in steps of hours to days) from the initial weather state only, using statistical relations learned from past data. Beyond improved forecasts, the major advantages of data-driven models are their substantially lower computational costs (and accompanied energy consumption) and the faster generation of forecasts, once a model has been trained. Over the past two years, fundamental advances have been achieved, with purely data-driven weather models now convincingly outperforming state-of-the-art NWP systems, as recently reviewed in Ben Bouallègue et al. (2024a). The most notable contributions and global models include Keisler (2022), FourCastNet (Pathak et al., 2022), Pangu-Weather (Bi et al., 2023), GraphCast (Lam et al., 2022), ClimaX (Nguyen et al., 2023a), FengWu (Chen et al., 2023a), FuXi (Chen et al., 2023c), SwinRDM (Chen et al., 2023b), AtmoRep (Lessig et al., 2023), NeuralGCM (Kochkov et al., 2023), Stormer (Nguyen et al., 2023b), GenCast (Price et al., 2023), and AIFS (Lang et al., 2024). All models utilize the ERA5 global reanalysis dataset (Hersbach et al., 2020) for training and evaluation, and are run at grid spacings of up to 0.25.
@@
However, most of these efforts have been focused on deterministic forecasts only, making it impossible to quantify forecast uncertainties which is crucial for optimal decision making, and one of the reasons underlying a transdisciplinary transition towards probabilistic forecasts (Gneiting and Katzfuss, 2014). Therefore, the overarching aim of our work is to investigate approaches to generate probabilistic predictions from deterministic data-driven weather models. An ideal solution to this challenge might be inherently probabilistic data-driven approaches, for example generative ML methods, and recent ensemble models such as AtmoRep (Lessig et al., 2023), NeuralGCM (Kochkov et al., 2023), GenCast (Price et al., 2023) or FuXi-ENS (Zhong et al., 2024) represent first steps in this direction. However, trained models are generally not yet publicly available and partly operate at different spatial resolutions than most of the deterministic data-driven weather models listed above.
By contrast, we consider readily applicable techniques to generate probabilistic forecasts. Specifically, we consider two main approaches for uncertainty quantification (UQ) for data-driven weather models. A schematic overview of these approaches is provided in Figure 1. Initial condition (IC)-based approaches generate an ensemble forecast by running a data-driven model multiple times based on a number of (slightly) different initial conditions. These initial condition ensembles can be generated in various ways, and we consider three variants: Adding random noise to the initial weather state, as proposed for example by Scher and Messori (2021) or Pathak et al. (2022); utilizing the perturbed initial conditions of a physics-based NWP ensemble model (Buizza et al., 2008); and generating conditions based on perturbations computed from randomly selected past data, as proposed by Magnusson et al. (2009). IC approaches generally require the capabilities to run the data-driven models for a set of input data (i.e., estimated models, code, data, and suitable hardware infrastructure), which has recently become possible since code and data have been made public for some of the models (most notably FourCastNet, Pangu-Weather, and GraphCast), but still poses technical challenges due to the substantial computing and disk space requirements.
@@
Post-hoc (PH) UQ approaches, by contrast, utilize statistical or ML methods to supplement deterministic forecasts with uncertainty information and thus turn them into probabilistic forecasts. These methods only require a training dataset of deterministic forecasts and corresponding observations. A large variety of such approaches has been proposed, e.g., conformal prediction (Angelopoulos and Bates, 2021) or distributional regression (Gneiting and Katzfuss, 2014). Here, we consider two distributional regression approaches particularly relevant for atmospheric science applications, where such methods have been mostly used in the context of statistical post-processing to correct systematic errors of NWP forecasts (Vannitsem et al., 2021). The EasyUQ (Walz et al., 2024a) approach builds on the recent isotonic distributional regression technique (IDR; Henzi et al., 2021) and yields statistically optimal discrete predictive distributions by leveraging the pool-adjacent-violators algorithm for nonparametric isotonic regression. EasyUQ utilizes deterministic forecasts of the target variable as sole input, and has, e.g., recently been used in Walz et al. (2024b) to generate probabilistic forecasts of precipitation from deterministic inputs. Over the past years, modern ensemble post-processing methods based on neural networks have been proposed which enable the incorporation of additional input variables and the data-driven learning of complex relationships between the inputs and distribution forecasts. We will build on the parametric distributional regression network approach first proposed in Rasp and Lerch (2018), which has been successfully extended for many target variables (e.g., in Schulz and Lerch, 2022) and has been used to generate corrected probabilistic forecasts from deterministic inputs from an NWP model (Chapman et al., 2022; Gneiting et al., 2023).
@@ 2 Data and setup
Our study focuses on the Pangu-Weather model developed by Bi et al. (2023). Additional results for the FourCastNet model (Pathak et al., 2022) are available in the supplemental material. Pangu-Weather is a three-dimensional vision transformer architecture with specific adaptations and extensions to weather prediction, and was one of the first data-driven models to achieve improvements over physics-based NWP models. The Pangu-Weather model produces global forecasts of five atmospheric variables (Z, Q, T, U, V) on 13 pressure levels and four surface variables (MSL, U10M, V10M, T2M) at a grid spacing of 0.25. It is trained based on 39 years of ERA5 reanalysis data from 1979–2017. In Bi et al. (2023), data from the year 2019 was used as validation data, and data from 2018 serves as a test dataset. For details regarding the model architecture, training procedure, and forecast quality, we refer to Bi et al. (2023). To implement the UQ methods described below, we adapted Pangu-Weather code and data provided by Bi et al. (2023) [^4] https://github.com/198808xc/Pangu-Weather for our purposes.
@@
As a reference forecast but also as initial condition perturbations (see Section 4), we retrieve operational ensemble forecasts of ECMWF’s ensemble prediction system, which is based on the ECMWF Integrated Forecasting System (IFS). The data are retrieved for the same training and evaluation periods on a regular latitude-longitude grid of 0.25$\times$0.25° covering the identical spatial domain and forecast lead times. It should be noted that the native spatial resolution of the operational ensemble prediction system of ECMWF is slightly higher than that of ERA5 which may cause differences in regions of high topography. For comparisons with post-processed IFS predictions in Section 55.2, we further utilize IFS forecast data available in WeatherBench 2 (Rasp et al., 2024).
3 Mathematical notation
@@ 4.1 Initial condition ensemble approaches
The general idea behind all considered initial condition ensemble approaches is that based on (slightly) different initial conditions
@@
The IC approaches described below mainly differ in the way the IC ensemble $\bm{X}_{l,t,0}$ is generated. Specifically, we consider Gaussian noise perturbations, random field perturbations and IFS perturbed initial conditions, which are introduced in detail below. Exemplary perturbations for a common initialization date are visualized in Figure 2. A natural shortcoming of all IC approaches is that they are inherently limited to accounting for initial condition uncertainty only, and not for model uncertainty, which is, e.g., addressed in physics-based NWP models via stochastic parametrizations of subgrid processes (Palmer, 2019b). One approach to address this has recently been proposed by MaheshEtAl2024, who construct IC perturbations with bred vectors and incorporate model uncertainty by utilizing an ensemble of data-driven weather models, the members of which have been trained separately from different random starting points.
@@ 4.1.1 Gaussian noise perturbations (GNP)
$X_{l,t,0}^{\text{GNP},m}=Y_{l,t}+\varepsilon_{l,t}^{m}\quad\text{for }m=1,...,M,$
where $\varepsilon_{l,t}^{m}\sim\mathcal{N}(0,\gamma\,\sigma)$, $\sigma$ denotes the mean standard deviation of the respective variable over all grid points, and $\gamma$ is a tuning parameter. Samples of the Gaussian noise process are thus generated independently over members $m$, locations $l$, variables and initialization times $t$. While Pathak et al. (2022) use $\gamma=0.3$ for the FourCastNet model, the Pangu-Weather model utilizes a substantially different architecture and models an increased number of meteorological variables. For our experiments we found that a scaling factor value of $\gamma=0.001$ applied to all variables works sufficiently well.
@@ 4.1.2 IFS initial conditions (IFSP)
Further, we consider an initial condition approach more akin to the operational practice of running NWP ensemble models (Palmer, 2019a) by utilizing the (ensemble of) initial conditions of the ECMWF ensemble prediction system to initialize the deterministic data-driven model. Specifically, we select the values at initialization time (i.e., the forecasts at step $h=0$) of the perturbed members $\bm{Z}_{t,0}^{m}=\{Z_{l,t,0}^{m},l\in\mathbb{L}_{G}\},m=1,...,M$, of the ECMWF ensemble, i.e.,
@@
4.1.3 Random field perturbations (RFP)
Finally, an alternative data-driven approach to generate IC ensembles, which we will refer to as random field perturbations, was proposed by Magnusson et al. (2009) in the context of physics-based NWP ensemble models. They argue that adding noise to the initial conditions ignores the underlying dynamics of the weather system. Instead, they suggest to use the scaled difference of two independent, randomly selected atmospheric states from the past as perturbation, which has the advantage of preserving linear balances in the system. The random field perturbations are calculated as
where $\bm{Y}_{\tau_{i}^{m}}=\{Y_{l,\tau_{i}^{m}},l\in\mathbb{L}_{G}\},i=1,2$ denotes the global observed ERA5 field of the selected variables at date $\tau_{i}^{m}$, and $\|\cdot\|_{\mathrm{Etot}}$ denotes the total energy norm, which is a conserved quantity of the governing equations of motion linearized about a reference state (cf. Magnusson et al., 2009). We choose the dates $\tau_{1}^{m},\tau_{2}^{m}$ randomly from the training dataset (2018–2021) and from the same month as $t$ to account for seasonal variability, but sample $\tau_{1}^{m}$ and $\tau_{2}^{m}$ from different years to ensure (approximate) independence. The constant $\alpha$ is a tuning parameter and controls the dispersion of the IC ensemble. Based on preliminary tests for a subset of initilization dates in which we tested the sensitivity of the spread-skill relationship to the magnitude of $\alpha$, we chose $\alpha=5\cdot 10^{6}$ for our case study. Increasing or decreasing the magnitude of $\alpha$ deteriorated the spread-skill relationship. Note that despite this scaling, the initial perturbation in terms of total energy are greater than with IFS initial conditions (Magnusson et al., 2009). With these choices, global perturbations $\bm{\xi}_{t}^{m,\alpha},m=1,...,M$, are computed and added to the corresponding ERA5 initial conditions, i.e.,
$X_{l,t,0}^{\text{RFP},m}=Y_{l,t}+\xi^{m,\alpha}_{l,t}$
for $m=1,...,M$ and all $l\in\mathbb{L}_{G}$.
@@ 4.2 Post-hoc approaches
is available, where $t$ denotes an initialization time in the training dataset (2018–2021).
Based on the training dataset, the PH methods yield forecast distributions $F_{l,t,h}$. Given a deterministic Pangu-Weather forecast $X^{\text{PW}}_{l,t^{\ast},h}$ in the test dataset (2022), a probabilistic forecast for the date $t^{\ast}$ and lead time $h$ can thus be obtained by using the corresponding deterministic forecast as input to the trained PH model. In the following, we consider two complementary post-hoc methods based on statistical and ML approaches.
@@ 4.2.1 EasyUQ
EasyUQ, proposed by Walz et al. (2024a), aims at learning a predictive distribution from deterministic, single-valued model output. As noted in the introduction, EasyUQ is a special case of IDR (Henzi et al., 2021) for a single deterministic prediction. EasyUQ proceeds separately for every location $l\in\mathbb{L}_{E}$ and lead time $h$. To simplify notation, we will suppress the lead time index $h$ in the current subsection, and note that all forecasts and observations should be understood as those for the corresponding lead time only. Given corresponding data of the form $(X^{\text{PW}}_{l,t},Y_{l,t}),t=1,...,T$, and assuming that the predictive cumulative distribution functions (CDFs) $F_{x}(y)=\mathbb{P}(Y_{l,t}\leq y|X_{l,t}=x)$ are increasing in stochastic order in $x$, i.e., $F_{x}(y)\geq F_{x^{\prime}}(y)$ for all $y\in\mathbb{R}$ if $x\leq x^{\prime}$, the EasyUQ-estimated predictive CDF is then given by
$\hat{F}^{\text{EasyUQ}}_{l,t}(y):=\hat{F}^{\text{EasyUQ}}_{X^{\text{PW}}_{l,t}}(y)=\min_{k=1,\ldots,t}\max_{\ell=t,\ldots,T}\frac{1}{\ell-k+1}\sum_{t^{\prime}=k}^{\ell}\mathbb{I}\{Y_{l,t^{\prime}}\leq y\},\quad t=1,\ldots,T.$
Thereby, $\hat{F}^{\text{EasyUQ}}_{l,t}$ is a statistically optimal discrete predictive distribution in that it minimizes the continuous ranked probability score (CRPS, see Section 44.3) over all conditional distributions satisfying the assumption of stochastic ordering. For theoretical results and more details on EasyUQ, we refer to Walz et al. (2024a) and Henzi et al. (2021). EasyUQ does not require any choices of tuning parameters and thus constitutes an attractive benchmark method that can be applied in a fully automated manner. Walz et al. (2024a) note that EasyUQ yields similar forecast performance as conformal prediction, and in case studies on post-processing, EasyUQ showed predictive performance comparable to other statistical methods (e.g., Schulz and Lerch, 2022).